OpenStudy (anonymous):

Choose the correct sum of the polynomials (4x3 - 2x - 9) + (2x3 + 5x + 3). 6x3 - 3x - 6 2x3 - 7x - 12 6x3 + 3x - 6 2x3 + 3x - 3 please help

OpenStudy (anonymous):

it's just addition. give it a try.

OpenStudy (anonymous):

\[\large 4x^3 - 2x - 9 + 2x^3 + 5x + 3 = ?\]

OpenStudy (anonymous):

can you walk me through it please

OpenStudy (anonymous):

Add the terms that have the same power together. \[\large 4x^3 + 2x^3 = ?\] \[\large 5x - 2x = ?\] \[\large 3 - 9 = ?\]

OpenStudy (anonymous):

What do you get from those?

OpenStudy (anonymous):

6x^3

OpenStudy (anonymous):

correct, keep going.

OpenStudy (anonymous):

3

OpenStudy (anonymous):

3x*

OpenStudy (anonymous):

and the last one is?

OpenStudy (anonymous):

lol yes that what i mean

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

Correct, so when you add them all back together you get: \[\large 6x^3 + 3x - 6\]

OpenStudy (anonymous):

thank you do you mind helping me with more

OpenStudy (anonymous):

sure.

OpenStudy (anonymous):

open up a new question.